OpenStudy (anonymous):

What is the center and radius of the circle with the equation x2 + y2 – 10x – 12y + 36 = 0 ?

OpenStudy (anonymous):

x2 – 10x + y2– 12y + 36 = 0 x2-10x+25+y2-12+36=-36+36+25 (x-5)^2+(y-6)^=25 were do i go from here ?

OpenStudy (anonymous):

gd! this is in the form of standard equation of a circle. (x-h)^2 + (y-k)^2 = (r)^2 compare your equation with this one. get the value of (h,k) which is co-ordinates of the centre of circle and r = radius/

OpenStudy (dan815):

yes put it in that form it will be fun >_>

OpenStudy (anonymous):

soo Center: (5, 6); how do i find the radius ?

OpenStudy (dan815):

complete the squarea and ull know the constant u will have left over

OpenStudy (dan815):

show ur work

OpenStudy (anonymous):

in your equation =25 is written =(5)^2 that means the radius = 5

OpenStudy (anonymous):

thankyou very much ! (:

OpenStudy (anonymous):

welcome!

OpenStudy (dan815):

oh u already solved it up there good job